# Floquet-engineered Emergent Massive Nambu–Goldstone Modes

### v3

@@ Abstract
Over the past decades, properties of mNGs have been extensively investigated in equilibrium [6, 7, 8, 9, 10, 11, 12, 13, 14, 15, 16, 17, 18, 19, 20, 21, 22, 23, 24, 25, 26, 27, 28, 29, 30, 31, 32, 33, 34, 35, 36], including, e.g., counting rules of their numbers [2], mass [11, 9] and damping processes [24, 28]. In contrast, little has been known in the case of generic time-dependent systems, which are ubiquitous in nature and experiments. Perhaps one of the simplest setups are time-periodic (Floquet) systems, which have been recently proposed as a toolbox to realize a zoo of novel non-equilibrium phases of matter [37, 38, 39, 40, 41, 42, 43, 44, 45, 46, 47, 48, 49, 50, 51, 52, 53, 54, 55]. Yet, the very existence of mNGs in Floquet systems remains largely unexplored, especially when symmetries are not exact. Here, we pose the questions of whether one can exploit Floquet protocols to manipulate various explicit SB processes [56, 57], stabilize mNGs, and even engineer intriguing non-equilibrium features without a counterpart in static systems.
@@
To address these challenges, we construct a Floquet protocol that leads to emergent mNGs with a parametrically tunable gap size and lifetime. The key conceptual ingredient is a driving scheme that dynamically exhibits an emergent hierarchical symmetry structure [56], together with a Lie group structure that guarantees the appearance of mNGs. The construction applies to classical and quantum systems alike, irrespective of the underlying microscopic details. Due to prethermalization-induced suppression of energy absorption at large drive frequencies (compared to local energy scales) [62, 63, 64], emergent mNGs can be identified well before the onset of heating. Moreover, we show that they respond differently to various types of explicit SB perturbations: their mass originates from effective processes preserving a subgroup structure, while their lifetime is determined by even weaker effects that eventually break all symmetries.
@@
Drive-induced higher-order processes break the $O(2)$ symmetry and destabilize the mNGs, leading to characteristic oscillations in quasi-conserved quantities, e.g., the $x/y/z$ components of the total magnetization. Their amplitudes exhibit distinctive scaling behavior, $A_{x/y}{\propto}T$ and $A_{z}{\propto}T^{2}$, confirming the hierarchical breakdown of the $O(3)$ symmetry in time. Despite their eventual heat death, these mNGs exhibit a parametrically long lifetime: the linewidth of the DSF scales as $\Gamma{\propto}T^{3}$ with a scaling exponent different from the naive Fermi’s Golden rule expectation. We justify it by applying a recently developed hydrodynamic theory designed to address the limit of weak explicit SB [24], as is the case in the high-frequency regime. As $\Gamma$ is much smaller than the gap size, drive-induced mNGs constitute well-defined quasi-particles of the Floquet system in the high-frequency regime.
@@
$U^{\pm}_{0}{=}e^{\pm iH_{0}T^{\prime}}$, $U_{1}^{\pm}{=}e^{\pm iH_{1}T^{\prime}/2}e^{\pm iH^{\prime}_{1}T^{\prime}/2}$, $U_{2}{=}e^{-iH_{2}T^{\prime}}$, where $T^{\prime}{=}{T}/{10}$ is the evolution time for each step and $T$ is the total drive period. To understand the resulting dynamics, one can define a static effective Hamiltonian $Q$ through the relation $U_{F}{\equiv}e^{-iQT}$. In the high-frequency regime, one can perturbatively determine $Q{=}\sum_{m}Q^{(m)}$, with $Q^{(m)}{\sim}T^{m}$, and its truncation at a finite order can be used to approximate the stroboscopic time evolution. Using an inverse-frequency expansion we obtain the two lowest order contributions
@@
and $H_{0}$ only appears in higher-order terms.
@@
In general, mNGs can be implemented by this protocol if $Q^{(0)}$ has a continuous symmetry. Yet, here we specifically require that $Q^{(0)}$ preserves one non-Abelian symmetry group, $G_{2}$, such that one can separately tune the mass and lifetime of mNGs by different higher-order processes, as elaborated below. For generic drives the first-order correction $Q^{(1)}$ breaks all subgroups of $G_{2}$. However, one can use the following construction to break the symmetry hierarchically. An illustrative case is when $H_{1}$ and $H^{\prime}_{1}$ are proportional to two non-commuting generators of $G_{2}$, say $X_{i}$ and $X_{j}$. Denoting the non-vanishing structure constants as $c_{ijk}$, with $[X_{i},X_{j}]{=}\sum_{k}c_{ijk}X_{k}$, allows us to write $Q^{(1)}$ as a superposition $Q^{(1)}{=}\mu\sum_{k}c_{ijk}X_{k}{\equiv}\mu M$ where $\mu$ denotes the chemical potential 111 The coefficients $c_{ijk}$ should contain non-vanishing elements otherwise the symmetry becomes Abelian. If so, the gap can only be generated by $Q^{(2)}$, which indeed also contains processes that destabilize mNGs; as a result, one cannot separately tune the gap and stability as in the non-Abelian case.. Consequently, $Q^{(0)}{+}Q^{(1)}$ preserves a subgroup $G_{1}{\subset}G_{2}$, spanned by the generators that commute with $M$. Suppose $|0\rangle$ labels the ground state of $Q^{(0)}{+}Q^{(1)}$ and the generator $M$ is broken (either explicitly or spontaneously). Ref. [66] has proved that generators that commute with $M$ can generate standard massless NGs, but non-commuting generators necessarily excite mNGs; the corresponding masses, exactly determined by group theory, are proportional to the chemical potential. In this example, $\mu$, and therefore the mass of the emergent mNGs, is linear in $T$.
@@
Since we do not assume any further symmetry structure in $H_{0}$, the symmetry group $G_{1}$ will be degraded to the trivial group by the next-order (in $T$) term $Q^{(2)}$. Therefore, although both explicit SB and SSB are allowed by the effective Hamiltonian $Q^{(0)}{+}Q^{(1)}$, only explicit SB occurs in the exact Floquet protocol. This is of significant importance when seeking quantitative descriptions of mNGs and two possible scenarios need to be addressed separately: |$\ket{0}$ is non-degenerate (Case 1) or degenerate (Case 2), w.r.t. the symmetry $G_{1}$.
@@
In Case 1, |$\ket{0}$ is unique and independent of $T$, and $Q^{(2)}$ slightly perturbs it by a negligible amount. Hence, $Q^{(0)}{+}Q^{(1)}$ is sufficient to predict the masses and dispersion of the mNGs to a good approximation. Case 2 is more subtle since $Q^{(2)}$ breaks $G_{1}$ and lifts the degeneracy; hence, one has to take at least $Q^{(2)}$ into account to determine the new ground state  [^2] In general, the new ground state will explicitly depend on $T$ and the concrete form of $Q^{(2)}$, which contains complicated nested commutators of each Hamiltonian in the drive, making the analytical prediction of the mNGs very difficult in practice. We also note that, the inverse-frequency expansion for Floquet systems does not converge in general. However, as a dynamic consequence of this divergence, Floquet heating can be delayed up to exponentially long times. We emphasize that this is irrelevant to the Goldstone physics discussed here, which only persists for algebraically long times.. For simplicity, in the following, we focus on Case 1 and briefly comment on Case 2 in the end; additional discussions are presented in the Supplementary Material (SM).
@@
Higher corrections of order $\mathcal{O}(T^{2})$ also lead to a finite lifetime of mNGs at any finite $T$; however, their influence can be systematically pushed to later times by increasing the driving frequency.
@@
The proposed construction is generic as it relies only on the aforementioned group structure; in particular, it is independent of the microscopic details of the drive, and hence can be applied to various physical systems of interest. The piece-wise constant protocol Eq. (1) can be implemented, e.g., on gate-based digital quantum simulators [60, 68]; in the SM we also present a continuous drive feasible for analog quantum platforms [69, 59, 70, 71].
@@
with $N$ sites and periodic boundary conditions. The normalized classical spin variable $\bm{S}_{j}$ on site $j$ can also be parametrized by azimuthal angle $\phi_{j}$ and polar angle $\theta_{j}$ as $S_{j}^{x}{=}\sin{\theta_{j}}\cos{\phi_{j}}$, $S_{j}^{x}{=}\sin{\theta_{j}}\sin{\phi_{j}}$, $S_{j}^{z}{=}\cos{\theta_{j}}$; it satisfies the Poisson bracket $\{S_{i}^{\mu},S_{j}^{\nu}\}{=}\delta_{ij}\epsilon^{\mu\nu\rho}S_{j}^{\rho}$ with the fully antisymmetric tensor $\epsilon^{\mu\nu\rho}$. Here $H_{2}$ corresponds to the Heisenberg model, and $H_{1},H_{1}^{\prime}$ are generators of $O(3)$ symmetries; $h_{x/y}$ are the magnetic field strength along the $x/y$-directions, and $J_{x}$ denotes the interaction strength along $x$-direction which breaks all continuous symmetries. The spin dynamics is described by $\dot{S}_{j}^{\mu}{=}\{S_{j}^{\mu},H(t)\}$, where the time-dependence in the Hamiltonian follows the piecewise protocol, Eq. (1). Following Eq. (2) we obtain the effective Hamiltonian
@@
with $q$ denoting the quasi-momentum, cf. SM for details.
@@
Figure 2: (a) Dynamical structure factor (DSF) through the first Brillouin zone. The left half displays numerical results, which precisely follow the theoretical prediction (right half) and certify the existence of mNGs. We use the drive period $T{=}2$. The Fourier transform is performed by using $10^{4}$ Floquet cycle. (b) A gap is opened at $q{=}0$ with a tunable size that depends linearly on $T$. We use $J_{2}{=}1$, $h_{x}{=}2$, $h_{y}{=}{-}2$, $J_{x}{=}1$, $\theta_{j}{=}0.01\pi$, and $\phi_{j}$ randomly sampled within $(-\pi/3,\pi/3)$.
@@
For each driving step, the time evolution is governed by the EOM generated by the corresponding static Hamiltonian, except for $H_{2}$ where a symmetric Trotter decomposition is used for numerical efficiency  [^4] Except for those generated by $H_{2}$, the EOMs can be explicitly integrated over the corresponding step duration. For $H_{2}$, we use a symmetric Trotter decomposition to approximate the evolution [77], thereby significantly improving the efficiency of the numerical simulations, cf. details in SM. This method introduces simulation errors at order $\mathcal{O}(T^{2})$, and hence it neither changes the form of $Q_{\mathrm{eff}}$, Eq. (4), nor gap size predicted above.. For all numerical results we use $N{=}1000$ and perform ensemble averages over $100$ realizations. The dynamics of the total magnetization components is shown in Fig. 1. The rotating magnetization in the $(x,y)$-plane (red and blue) is a characteristic signature of the mNGs. The oscillation is periodic in time (clearly visible in the inset where a linear time scale is used) and persists for a controllably long time without exhibiting any notable decay up to $\ell T{\approx}10^{5}.$ Although $Q_{\text{eff}}$ in Eq. (4) conserves the $z$-magnetization and predicts a static $\langle S_{z}^{\text{tot}}\rangle/N$ (dashed line), the exact Floquet protocol indeed generates an oscillatory dynamics (green) since $Q^{(2)}$ breaks the $O(2)$ symmetry. Oscillations in the $(x,y)$-plane gradually damp out and settle into a prethermal plateau, which eventually heat up to an infinite temperature ensemble, cf. SM.
@@
$Q^{(2)}$ that cannot be captured by Eq. (4) breaks all symmetries, and hence modify and destabilize the mNGs. The most notable phenomenon is the oscillatory dynamics in $\langle S_{z}^{\text{tot}}\rangle/N$ as seen before in Fig. 1: it occurs because higher order corrections can effectively generate extra fields in the $(x,y)$-plane, slightly tilting the original explicit SB generator of the $O(2)$ symmetry in $Q_{\mathrm{eff}}$. For low-temperature initial states and at short times, the three spin directions are indeed all coupled together by this field; therefore, they share the same oscillating frequency, cf. inset of Fig. 1. However, one clearly observes a difference in the oscillating amplitudes and $A_{x/y}$ is significantly larger than $A_{z}$. More precisely, in Fig. 3 we confirm their asymptotic scaling dependence on $T$, $A_{x/y}{\sim}T$ and ${A_{z}\sim}T^{2}$, which can also be analytically justified by a perturbative argument, cf. SM.
@@
For even longer times, higher order effects together with thermal fluctuations (set by the spatial randomness in the initial states) lead to the damping of mNGs which broadens the spectrum. In Fig. 4 (a) we depict the Fourier spectrum for $q{=}0$. Strikingly, we observe an asymmetric lineshape and multiple side peaks of comparable heights, one intriguing non-equilibrium feature that normally does not exist in systems without driving. They depend on the details of the Trotterization protocol being used and also on the specific form of the SB processes, cf. SM.
The overall linewidth $\Gamma$ in the DSF can be used to characterize the damping rate. It is determined by averaging the full width at $10\%$, $15\%$, and $20\%$ of the maximum intensity of the Fourier peak, with their standard deviation serving as the error bar. $\Gamma$ exhibits a power-law scaling $T^{\alpha}$, suggesting that the mNGs lifetime can be systematically prolonged by increasing the drive frequency. Since all symmetries are broken by $\mathcal{O}(T^{2})$ processes, following Fermi’s Golden rule one may expect a scaling exponent $\alpha{=}4$ [80, 81, 82]; however, surprisingly, we find $\alpha{\approx}3$, significantly deviating from this naive expectation.
@@
Discussion.— The scaling exponent of the mNGs lifetime may not be universal and can depend on Hamiltonian parameters. For instance, for the orange data shown in Fig. 4(b) with substantially stronger $O(2)$ SB effects, $\alpha$ is indeed larger than 3, see more examples in SM. To explain this scaling behavior one needs to further analyse the scattering channels between mNGs modes, e.g. the two- or three-magnon scattering matrix. Given different SB processes and thermal fluctuations, the dominant effects in $D_{0}$ may scale differently with $T$, and will be systematically studied in future work. As the scattering matrix now becomes tunable via Floquet-engineering, our work paves a new pathway to stabilize or create novel magnon bound states  [58, 86, 61, 87, 88], as well as different transport behaviors which may not exist in non-driven systems.
While we have discussed the emergent mNGs in a classical Floquet system, we emphasize that our Floquet protocol applies equally to both quantum and classical systems. In SM, we present numerical simulations of a quantum system and certify the existence of mNGs [89], illustrating the possibility for experimentally detecting the mNGs in real quantum simulators.
Going beyond the present systems, it is intriguing to apply our general construction to physical systems exhibiting different types of mNGs. For instance, one can consider the Heisenberg model with anti-ferromagnetic couplings: a weak external $z$-field leads to a ground state manifold which preserves a $O(2)$ symmetry but this degeneracy is then lifted by $Q^{(2)}$ in driven systems (Case 2). As illustrated in SM, one mNG can still be detected from the spin dynamics. However, we observe strong evidence that the damping effect is substantially different from the FM case, opening up new future directions  [^5] Indeed, $Q^{(0)}+Q^{(1)}$ also predicts a gapless mode with linear dispersion. However, it is unstable against higher order perturbations and we do not observe them in the exact Floquet dynamics..
Furthermore, the requirement that $Q^{(1)}$ forms a superposition of generators of the non-Abelian group $G_{2}$, is not necessary for the appearance of mNGs, so long as $Q^{(1)}$ still preserves a subsymmetry of $G_{2}$. We give one explicit example in the SM where $Q^{(1)}$ involves three-body interactions, which effectively generates a $z$-field for sufficiently small thermal fluctuations in the initial state. Yet, unlike the model in Eq. (4) where the gap size can be exactly determined, here the gap can only be computed approximately. Also, this mode tends to be very fragile with a notable broadening of the spectrum linewidth that scales linearly in $T$, see details in SM.
@@
Acknowledgments.— This work is supported by “The Fundamental Research Funds for the Central Universities, Peking University”, and by ”High-performance Computing Platform of Peking University” and by the Deutsche Forschungsgemeinschaft under the cluster of excellence ct.qmat (EXC 2147, project-id 390858490). This work is funded by the European Union (ERC, QuSimCtrl, 101113633). Views and opinions expressed are however those of the authors only and do not necessarily reflect those of the European Union or the European Research Council Executive Agency. Neither the European Union nor the granting authority can be held responsible for them.

### v4

@@ Abstract
Over the past decades, properties of mNGs have been extensively investigated in equilibrium [6, 7, 8, 9, 10, 11, 12, 13, 14, 15, 16, 17, 18, 19, 20, 21, 22, 23, 24, 25, 26, 27, 28, 29, 30, 31, 32, 33, 34, 35, 36], including, e.g., counting rules for their numbers [2], mass [11, 9] and damping processes [24, 28]. In contrast, little has been known in the case of generic time-dependent systems, which are ubiquitous in nature and experiments. Perhaps one of the simplest setups are time-periodic (Floquet) systems, which have been recently proposed as a toolbox to realize a zoo of novel non-equilibrium phases of matter [37, 38, 39, 40, 41, 42, 43, 44, 45, 46, 47, 48, 49, 50, 51, 52, 53, 54, 55, 56, 57]. Yet, the very existence of mNGs in Floquet systems remains largely unexplored, especially when symmetries are not exact. Here, we pose the questions of whether one can exploit Floquet protocols to manipulate various explicit SB processes [58, 59], stabilize mNGs, and even engineer intriguing non-equilibrium features without a counterpart in static systems.
@@
To address these challenges, we construct a family of Floquet protocols that lead to emergent mNGs with a parametrically tunable gap size and lifetime. The key conceptual ingredient is a driving scheme that dynamically exhibits an emergent hierarchical symmetry structure [58], together with a Lie group structure that guarantees the appearance of mNGs. The construction applies to classical and quantum systems alike, irrespective of the underlying microscopic details. Due to prethermalization-induced suppression of energy absorption at large drive frequencies (compared to local energy scales) [64, 65, 66], emergent mNGs can be identified well before the onset of heating. Moreover, we show that they respond differently to various types of explicit SB perturbations: their mass originates from effective processes preserving a subgroup structure, while their lifetime is determined by even weaker effects that eventually break all symmetries.
@@
Drive-induced higher-order processes break the $O(2)$ symmetry and destabilize the mNGs, leading to characteristic oscillations in quasi-conserved quantities, e.g., the $x/y/z$ components of the total magnetization. Their amplitudes exhibit distinctive scaling behavior, $A_{x/y}{\propto}T$ and $A_{z}{\propto}T^{2}$, confirming the hierarchical breakdown of the $O(3)$ symmetry in time. Despite their eventual heat death, these mNGs exhibit a parametrically long lifetime: the linewidth of the DSF scales as $\Gamma{\propto}T^{3}$ with a scaling exponent different from the naive Fermi’s Golden rule expectation. We rationalize this scaling by applying a recently developed hydrodynamic theory designed to address the limit of weak explicit SB [24], as is the case in the high-frequency regime. As $\Gamma$ is much smaller than the gap size, drive-induced mNGs constitute well-defined quasi-particles of the Floquet system in the high-frequency regime.
@@
$U^{\pm}_{0}{=}e^{\pm iH_{0}T^{\prime}}$, $U_{1}^{\pm}{=}e^{\pm iH_{1}T^{\prime}/2}e^{\pm iH^{\prime}_{1}T^{\prime}/2}$, $U_{2}{=}e^{-iH_{2}T^{\prime}}$, where $T^{\prime}{=}{T}/{10}$ is the evolution time for each step and $T$ is the total drive period. To understand the resulting dynamics, one can define a static effective Hamiltonian $Q$ through the relation $U_{F}{\equiv}e^{-iQT}$. In the high-frequency regime, one can perturbatively determine $Q{=}\sum_{m}Q^{(m)}$, with $Q^{(m)}{\sim}T^{m}$, and its truncation at a finite order, $Q_{\mathrm{eff}}$, can be used to approximate the stroboscopic time evolution. Using an inverse-frequency expansion we obtain the two lowest order contributions
@@
and $H_{0}$ only appears in higher-order terms. At early times, the system evolves into a long-lived prethermal plateau with quasi-conserved energy. This prethermal state can be captured by the Gibbs ensemble $\exp(-\beta Q_{\mathrm{eff}})$ [67], with the inverse temperature $\beta$ determined by the initial state energy  [^1] In this letter, temperature is denoted by $1/\beta$ to avoid confusion with the driving period $T$.. At this stage, concepts and methods in equilibrium systems remain valid, which we use in the following to predict characteristic properties of mNGs and interpret the collective motion of our driven system.
@@
In general, mNGs can be implemented by this protocol if $Q^{(0)}$ has a continuous symmetry. Yet, here we specifically require that $Q^{(0)}$ preserves one non-Abelian symmetry group, $G_{2}$, such that one can separately tune the mass and lifetime of mNGs by different higher-order processes, as elaborated below. For generic drives the first-order correction $Q^{(1)}$ breaks all subgroups of $G_{2}$. However, one can use the following construction to break the symmetry hierarchically. An illustrative case is when $H_{1}$ and $H^{\prime}_{1}$ are proportional to two non-commuting generators of $G_{2}$, say $X_{i}$ and $X_{j}$. Denoting the non-vanishing structure constants as $c_{ijk}$, with $[X_{i},X_{j}]{=}\sum_{k}c_{ijk}X_{k}$, allows us to write $Q^{(1)}$ as a superposition $Q^{(1)}{=}\mu\sum_{k}c_{ijk}X_{k}{\equiv}\mu M$ where $\mu$ denotes the chemical potential  [^2] The coefficients $c_{ijk}$ should contain non-vanishing elements, otherwise the symmetry becomes abelian. Then $Q^{(2)}$ breaks the remaining symmetry and opens the mNG gap. However, due to the absence of the non-abelian group structure, determining this gap becomes complicated and requires a case-by-case study. Furthermore, since $Q^{(2)}$ does not preserve any symmetry, mNGs can be quickly destabilized, see detailed discussions in the Sec. SM 4 [71]..
Consequently, $Q^{(0)}+Q^{(1)}$ preserves a subgroup $G_{1}\subset G_{2}$, spanned by the generators that commute with $M$. Suppose |0|0$\rangle$| 0 ⟩ labels the ground state of $Q^{(0)}{+}Q^{(1)}$ and the generator $M$ is broken (either explicitly or spontaneously). Ref. [9] showed that generators that commute with $M$ can generate standard massless NGs, but non-commuting generators necessarily excite mNGs; the corresponding masses, exactly determined by group theory, are proportional to the chemical potential. In this example, $\mu$, and therefore the mass of the emergent mNGs, is linear in $T$.
@@
Since we do not assume any further symmetry structure in $H_{0}$, the symmetry group $G_{1}$ will be degraded to the trivial group by the next-order (in $T$) term $Q^{(2)}$. Therefore, although both explicit SB and SSB are allowed by the effective Hamiltonian $Q^{(0)}{+}Q^{(1)}$, only explicit SB occurs in the exact Floquet protocol. This is of significant importance when seeking quantitative descriptions of mNGs, and two possible scenarios need to be addressed separately: |$\ket{0}$ is non-degenerate (Case 1) or degenerate (Case 2), w.r.t. the symmetry $G_{1}$.
@@
In Case 1, |$\ket{0}$ is unique and independent of $T$, and $Q^{(2)}$ is a negligible perturbation. Hence, $Q^{(0)}{+}Q^{(1)}$ is sufficient to predict the masses and dispersion of the mNGs to a good approximation. Case 2 is more subtle since $Q^{(2)}$ breaks $G_{1}$ and lifts the degeneracy; hence, one has to take at least $Q^{(2)}$ into account to determine the new ground state  [^3] In general, the new ground state will explicitly depend on $T$ and the concrete form of $Q^{(2)}$, which contains complicated nested commutators of each Hamiltonian in the drive, making the analytical prediction of the mNGs very difficult in practice. We also note that the inverse-frequency expansion for Floquet systems does not converge in general. However, as a dynamic consequence of this divergence, Floquet heating can be delayed up to exponentially long times. We emphasize that this is irrelevant to the Goldstone physics discussed here, which only persists for algebraically long times.. For simplicity, in the following, we focus on Case 1 and briefly comment on Case 2 in the end; additional discussions are presented in the Supplementary Material (SM) [71].
@@
Higher corrections of order $\mathcal{O}(T^{2})$ also lead to a finite lifetime of mNGs at any finite drive period $T$; however, their influence can be systematically pushed to later times by increasing the driving frequency.
@@
The proposed construction is generic as it relies only on the aforementioned group structure; in particular, it is independent of the microscopic details of the drive, and hence can be applied to various physical systems of interest. Indeed, we emphasize that there exist many other protocols to Floquet engineer mNGs with similar properties. In the Sec. SM 1.1 [71], we provide a comprehensive overview of a family of protocols (some of them involving fewer steps than Eq. (1)), which can be efficiently implemented, e.g., on gate-based digital quantum simulators [50, 62, 72]; in Sec. SM 1.2 [71] we also present a continuous drive feasible for analog quantum platforms [73, 61, 74, 75]. This variety of different protocols also highlights that our proposed Floquet-engineered mNGs is a fundamental and fairly universal emergent phenomenon.
@@
with $N$ sites and periodic boundary conditions. The normalized classical spin variable $\bm{S}_{j}$ on site $j$ can also be parametrized by azimuthal angle $\phi_{j}$ and polar angle $\theta_{j}$ as $S_{j}^{x}{=}\sin{\theta_{j}}\cos{\phi_{j}}$, $S_{j}^{y}{=}\sin{\theta_{j}}\sin{\phi_{j}}$, $S_{j}^{z}{=}\cos{\theta_{j}}$; it satisfies the Poisson bracket $\{S_{i}^{\mu},S_{j}^{\nu}\}{=}\delta_{ij}\epsilon^{\mu\nu\rho}S_{j}^{\rho}$ with the fully antisymmetric tensor $\epsilon^{\mu\nu\rho}$. Here $H_{2}$ corresponds to the Heisenberg model, and $H_{1},H_{1}^{\prime}$ are generators of $O(3)$ symmetries;
$h_{x/y}$ are the magnetic field strength along the $x/y$-directions, and $J_{x}$ denotes the interaction strength along $x$-direction which breaks all continuous symmetries. The spin dynamics is described by $\dot{S}_{j}^{\mu}{=}\{S_{j}^{\mu},H(t)\}$, where the time-dependence in the Hamiltonian follows the piecewise protocol, Eq. (1). Following Eq. (2) we obtain the effective Hamiltonian
@@
with $q$ denoting the quasi-momentum, cf. Sec. SM 2.1 [71] for details.
@@
Figure 2: (a) Dynamical structure factor (DSF) through the first Brillouin zone. The left half displays numerical results, which precisely follow the theoretical prediction (right half) and certify the existence of mNGs. We use the drive period $T{=}2$. The Fourier transform is performed by using $10^{4}$ Floquet cycles. (b) A gap is opened at $q{=}0$ with a tunable size that depends linearly on $T$. The gap size deviates from the theoretical prediction at large $T$. We use $J_{2}{=}1$, $h_{x}{=}2$, $h_{y}{=}{-}2$, $J_{x}{=}1$, $\theta_{j}{=}0.01\pi$, and $\phi_{j}$ randomly sampled within $(-\pi/3,\pi/3)$.
@@
We further verify the entire mNG spectrum by analyzing the DSF $\mathcal{S}(q,\omega){=}\sum_{\alpha{=}x,y}\langle|{S}_{q}^{\alpha}(\omega)|\rangle$, where ${S}_{q}^{\alpha}(\omega)$ is the space-time Fourier transform of the spin configurations, ${S}_{q}^{\alpha}(\omega){=}\sum_{j}\int\text{d}t{S}_{j}^{\alpha}(t)e^{-i(\omega t-qj)}$. Fig. 2(a) depicts $\mathcal{S}(q,\omega)$ for a fixed $T$: the spectrum precisely follows our theoretical prediction (green line in the right half), Eq. (5). A mNG gap opens at $q{=}0$ corresponding to the oscillatory frequency of the total magnetization. It is worth highlighting the controllability of the gap size by tuning the drive period, which also exhibits quantitative agreement with the theoretical prediction (green dashed line) over a wide range of $T$. A noticeable deviation appears when $T$ becomes large, $T\geq 3.5$, as shown in Fig. 2(b). Note, a faint gapless line also appears in Fig. 2(a), which is generated by non-linear and drive-induced higher-order effects, see details in Sec. SM 2.3 [71].
$Q^{(2)}$ that is not captured by Eq. (4) breaks all symmetries, and hence modifies and destabilizes the mNGs. The most notable phenomenon is the oscillatory dynamics in $\langle S_{z}^{\text{tot}}\rangle/N$ as seen before in Fig. 1: it occurs because higher order corrections can effectively generate extra fields in the $(x,y)$-plane, slightly tilting the original explicit SB generator of the $O(2)$ symmetry in $Q_{\mathrm{eff}}$. For low-temperature initial states and at short times, the three spin directions are indeed all coupled together by this field; therefore, they share the same oscillating frequency, cf. inset of Fig. 1. However, one clearly observes a difference in the oscillating amplitudes, with $A_{x/y}$ significantly larger than $A_{z}$. More precisely, in Fig. 3 we confirm their asymptotic scaling dependence on $T$, $A_{x/y}{\sim}T$ and ${A_{z}\sim}T^{2}$, which can also be analytically justified by a perturbative argument, cf. Sec. SM 2.2 [71].
For even longer times, higher order effects together with thermal fluctuations (set by the spatial randomness in the initial states) lead to the damping of mNGs which broadens the spectrum. In Fig. 4 (a) we depict the Fourier spectrum at $q{=}0$. Strikingly, we observe an asymmetric lineshape and multiple side peaks of comparable heights, one intriguing non-equilibrium feature that normally does not exist in systems without driving. They depend on the details of the Trotterization protocol being used and also on the specific form of the SB processes, cf. Sec. SM 2.4 [71].
[…]
@@
Discussion.— The scaling exponent of the mNGs lifetime may not be universal and can depend on Hamiltonian parameters. For instance, for the orange data shown in Fig. 4(b) with substantially stronger $O(2)$ SB effects, $\alpha$ is indeed larger than 3, see more examples in Sec. SM 2.6 [71]. To explain this scaling behavior one needs to further analyse the scattering channels between mNGs modes, e.g. the two- or three-magnon scattering matrix. Given different SB processes and thermal fluctuations, the dominant effects in $D_{0}$ may scale differently with $T$, and will be systematically studied in future work. As the scattering matrix now becomes tunable via Floquet-engineering, our work paves a new pathway to stabilize or create novel magnon bound states  [60, 90, 63, 91, 92], as well as different transport behaviors which may not exist in non-driven systems. Furthermore, for other values of $q$, $\Gamma$ scales differently with $T$; in particular, $\Gamma$ becomes almost independent of $T$ for large $q$, indicating that thermal fluctuations in the initial state and $Q^{(0)}$ dominate the damping, see details in Sec. SM 2.5 [71].
While we have discussed the emergent mNGs in a classical Floquet system, we emphasize that our Floquet protocol applies equally to both quantum and classical systems. In Sec. SM 2.9 [71], we present numerical simulations of a quantum system and certify the existence of mNGs [93], illustrating the possibility for experimentally detecting the mNGs in real quantum simulators.
Going beyond the present systems, it is intriguing to apply our general construction to physical systems exhibiting different types of mNGs. For instance, one can consider the Heisenberg model with anti-ferromagnetic couplings: a weak external $z$-field leads to a ground state manifold which preserves a $O(2)$ symmetry but this degeneracy is then lifted by $Q^{(2)}$ in driven systems (Case 2). As illustrated in Sec. SM 3 [71], one mNG can still be detected from the spin dynamics. However, we observe strong evidence that the damping effect is substantially different from the FM case, opening up new future directions  [^6] Indeed, $Q^{(0)}+Q^{(1)}$ also predicts a gapless mode with linear dispersion. However, it is unstable against higher order perturbations and we do not observe them in the exact Floquet dynamics..
Furthermore, the requirement that $Q^{(1)}$ forms a superposition of generators of the non-Abelian group $G_{2}$, is not necessary for the appearance of mNGs, so long as $Q^{(1)}$ still preserves a subsymmetry of $G_{2}$. We give one explicit example in the Sec. SM 2.8 [71] where $Q^{(1)}$ involves three-body interactions, which effectively generates a $z$-field for sufficiently small thermal fluctuations in the initial state. Yet, unlike the model in Eq. (4) where the gap size can be exactly determined, here the gap can only be computed approximately. Also, this mode tends to be very fragile with a notable broadening of the spectrum linewidth that scales linearly in $T$, see details in Sec. SM 2.8 [71].
@@
Acknowledgments.—This work is supported by the National Natural Science Foundation of China (Grant No. 12474214), and by Innovation Program for Quantum Science and Technology (No. 2024ZD0301800), and by Beijing Natural Science Foundation (Grant No. QY24016) and by the Deutsche Forschungsgemeinschaft under the cluster of excellence ct.qmat (EXC 2147, project-id 390858490), and by “The Fundamental Research Funds for the Central Universities, Peking University”, and by ”High-performance Computing Platform of Peking University”. This work is funded by the European Union (ERC, QuSimCtrl, 101113633). Views and opinions expressed are however those of the authors only and do not necessarily reflect those of the European Union or the European Research Council Executive Agency. Neither the European Union nor the granting authority can be held responsible for them.